# The effect of accretion on scalar superradiant instability

### v1

@@ IV Multi-mode evolution with accretion
In this section, we study the evolution in the presence of multiple modes. The evolution can be separated into stages with different $l$. In each $l$-stage, the modes with the same azimuthal number $l$ grow the fastest via superradiance. We keep only the dominant ($\{0,l,l\}$) and the subdominant ($\{1,l,l\}$) modes in each $l$-stage for illustration, while the generalization to more modes is straightforward. The superradiance rates of modes with azimuthal numbers larger than $l$ are orders of magnitude smaller and thus can be safely ignored in the $l$-stage. From the experience learned in the previous section, we start with the BH evolution on the Regge plot in Sec. IV.1. Readers not interested in technical details can only read this subsection. Analytical approximations for the important quantities are given in Sec. IV.2. Then the subdominant modes and GW emission flux are discussed in Sec. IV.3 and Sec. IV.4, respectively.
IV.1 Evolution on the Regge plot
@@
We first analyze the case when the initial BH spin $a_{*0}$ is greater than the $a_{*\text{c}}^{(0l_{\mathrm{c}}l_{\mathrm{c}})}$ given in Eq. (10). If the accretion rate is much larger than the superradiance rate of the $l_{c}$ mode, the BH first follows the accretion trajectory to $\sim a_{*\text{lim}}$ before transiting to the Regge curve with $l=l_{\text{c}}$, see Fig. 6(a). With a decreasing value of the accretion rate, the time cost on the accretion trajectory is shorter, see Fig. 6(b). At some point, the BH spin converges to the $l_{\text{c}}$ Regge curve since $t=0$, see Fig. 6(c). Further decreasing the accretion rate and/or increasing the superradiance rate, the time cost along the $l_{c}$ Regge trajectory is shorter, see Fig. 6(g,h). Until at some point, the accretion effect is negligible during the time when the BH is on the $l_{\text{c}}$ Regge curve. Then the BH does not evolve along the $l_{\text{c}}$ Regge curve before stepping down to the $l_{\text{c}}+1$ Regge trajectory, see Fig. 6(i).
Given the initial conditions, we need to judge which pattern the later evolution has. To separate these cases, one first calculates $\alpha_{0^{\prime}}^{(0l_{\text{c}}l_{\text{c}})}$, which is proportional to the BH mass at the point when its spin reaches the local maximum close to the trajectory curve. The formula for $l_{\text{c}}=1$ is given in Eq. (58), and those for larger values of $l_{\text{c}}$ are listed in Eqs. (90). When $\mu\tau_{\text{acc}}$ is large, the asymptotic expression of the Lambert $W_{0}$ function could be useful
$\lim_{x\to+\infty}W_{0}(x)=\log x-\log\left(\log x\right).$ (62)
@@
In any of the above three cases, the evolution merges onto the $\{0,l_{\text{c}},l_{\text{c}}\}$ Regge curve. With only a single mode as explained in Sec. III, the BH spin reaches a local minimum close to the Regge curve. There the time of the local minimum is defined as $t_{1}$, and we also define the quantities at this time using the subscript, such as $\alpha_{1}$. In the presence of multiple modes, there is a local minimum of the BH spin at each Regge curve. We similarly define this time at the $\{0,l,l\}$ Regge curve as $t_{1}^{(0ll)}$ and the corresponding quantities using an additional superscript $(0ll)$, such as $\alpha_{1}^{(0ll)}$. To avoid confusion, the mass of a superradiant mode at a certain time is written explicitly. For example $M_{\text{s}}^{(022)}(t_{1}^{(011)})$ denotes the mass of the $\{0,2,2\}$ mode at time $t_{1}^{(011)}$.
The value of $\alpha_{1}^{(0l_{\text{c}}l_{\text{c}})}$ can be obtained using Eqs. (91). One could then calculate the value of $\alpha_{1}^{(0ll)}$ with $l>l_{\text{c}}$ using Eqs. (103). Then the evolution on the Regge plot can be approximately fixed with the values of $\alpha_{0^{\prime}}$ and $\alpha_{1}^{(0ll)}$. When $\mu\tau_{\text{acc}}$ is large, the following asymptotic expression for the Lambert $W_{-1}$ function is useful
$\lim_{x\to 0^{-}}W_{-1}(x)=\log(-x)-\log\left(-\log(-x)\right).$ (63)
@@
$\begin{split}&{a}_{\mathrm{*c}}^{(0,l+1,l+1)}\left(\alpha^{(0,l+1,l+1)}_{1}\right)^{2}+(l+1)\mu{M}_{\mathrm{s}}^{(0,l+1,l+1)}(t_{1}^{(0ll)})\\ &\hskip 28.45274pt={a}_{\mathrm{*c}}^{(0ll)}\left(\alpha^{(0ll)}_{1}\right)^{2}+l\mu{M}_{\mathrm{s}}^{(0ll)}(t_{1}^{(0ll)}),\end{split}$ (64a)
$\alpha_{1}^{(0,l+1,l+1)}+\mu{M}_{\mathrm{s}}^{(0,1+1,1+1)}(t_{1}^{(0ll)})=\alpha_{1}^{(0ll)}+\mu{M}_{\mathrm{s}}^{(0ll)}(t_{1}^{(0ll)}).$ (64b)
The $\alpha_{1}^{(0ll)}$ obtained from these conservation equations agrees better with the numerical results, which can be observed in Fig. 6(c,g-i). Since the conservations give $\alpha_{1}^{(0,l+1,l+1)}\approx 0.9\,\alpha_{1}^{(0ll)}$ Guo:2022mpr , we claim Eqs. (64) give a more accurate description when $\alpha_{1}^{(0,l+1,l+1)}-\alpha_{1}^{(0ll)}\ll 0.1\,\alpha_{1}^{(0ll)}$.
@@
$a_{*\text{c}}^{(0,l^{\prime}+1,l^{\prime}+1)}<a_{*0}<a_{*\text{c}}^{(0,l^{\prime},l^{\prime})}.$ (65)
@@
We define $l_{\text{acc}}$ as the first Regge curve which the evolution follows along. Its value could be larger, smaller, or equal to $l^{\prime}$, depending on the relative size of the accretion rate and the superradiance rates of different modes. We first judge whether $l_{\text{acc}}$ is larger than $l^{\prime}$. After many tests, the best scheme we could find to calculate $l_{\text{acc}}$ is by comparing the $\alpha_{0^{\prime}}^{(0ll)}$ with different $l$. Here $\alpha_{0^{\prime}}^{(0ll)}$ is the mass coupling at the time when the BH spin reaches its local maximum if only accretion and the $\{0,l,l\}$ mode are present, which is given by Eqs. (90). Specifically, one calculate the $\alpha_{0^{\prime}}^{(0ll)}$ with increasing value of $l$ from $l_{\text{c}}$ to $l^{\prime}$. If a minimum of $\alpha_{0^{\prime}}^{(0ll)}$ is observed for $l<l^{\prime}$, then the BH spin reaches the local maximum between the $\{0,l-1,l-1\}$ and $\{0,l,l\}$ Regge curves. In this case, we set $l_{\text{acc}}=l$. After identifying $l_{\text{acc}}$, one could reset $t=0$ as the time when the BH reaches the intersection of the accretion trajectory and the $\{0,l_{\text{acc}},l_{\text{acc}}\}$ Regge curve. Then the later evolution can be described in the same way as the $a_{*0}>a_{*\text{c}}^{(0l_{\text{c}}l_{\text{c}})}$ case with $l_{\text{c}}$ replaced by $l_{\text{acc}}$.
@@
The evolutions with $a_{*0}=0$ are presented in Fig. 6, panels (d-f) and panels (j-l), where other initial parameters are the same as panels (a-c) and panels (g-i), respectively. In this case, the $l^{\prime}$ is infinity. Analytical estimates of the points where the evolution leaves a trajectory and merges into the next are given in Sec. IV.2. For various initial conditions and accretion rates, the analytical approximations agree quite well with the numerical results.
IV.2 Evolution of the BH and dominate modes
@@
The BH evolution is mainly determined by the dominant mode and/or the accretion in each $l$-stage. In this part, we focus on the dominant modes, while leaving the discussion about the subdominant modes in Sec. IV.3. To help the analysis, we present the numerical results for a specific multi-mode scenario in Fig. 7. The initial parameters are $M_{0}=10^{4}M_{\odot}$, $\alpha_{0}=0.1$, $M_{\text{s,0}}^{(nll)}=10^{-5}M_{0}$ and $f_{\text{Edd}}=10^{-3}$. The value of $f_{\text{Edd}}$ is chosen such that the accretion time scale ${\tau}_{\text{acc}}$ is at the same order of $\tau_{\text{SR}}^{(022)}$, which is the the superradiance time scale of $(0,2,2)$ mode. This time scale is much larger than $\tau_{\text{SR}}^{(011)}$ but much smaller than $\tau_{\text{SR}}^{(033)}$.
@@
Two initial BH spins $a_{*0}=0.7$ and $a_{*0}=0$ are considered. From Eq. (61), it is easy to see $l_{\text{c}}=1$. Then $a_{*\text{c}}^{(011)}\approx 0.384$ is calculated from Eq. (10) using the value of $\alpha_{0}$.
IV.2.1 $a_{*0}=0.7$
@@
This case corresponds to $a_{*0}>a_{*\text{c}}^{(0l_{\mathrm{c}}l_{\mathrm{c}})}$. We first identify the evolution of the BH on the Regge plot. From Eq. (18), one could obtain $\tau_{\text{acc}}\mu=3.2\times 10^{18}$. We then apply Eq. (58) and get $\alpha_{0^{\prime}}=\alpha_{0}-8\times 10^{-10}$, which is less than $\alpha_{0}$. Thus the BH does not evolve along the accretion trajectory at all, corresponding to the third scenario in the discussion below Eq. (62).
@@
The $\alpha_{1}^{(011)}$ can be calculated using Eq. (44). Since the BH does not evolve along the accretion trajectory, the $M_{\text{s},0^{\prime\prime}}$ $a_{*0^{\prime\prime}}$, and $\alpha_{0^{\prime\prime}}$ should be replaced by $M_{\text{s,0}}$, $a_{*0}$, and $\alpha_{0}$, respectively. Using the above parameters, one obtains $\alpha_{1}^{(011)}=\alpha_{0}+7\times 10^{-10}$. The small values of $\alpha_{0^{\prime}}-\alpha_{0}$ and $\alpha_{1}^{(011)}-\alpha_{0}$ indicate the evolution merges onto the $\{0,1,1\}$ Regge curve almost at $t=0$.
@@
We continue to calculate $\alpha_{1}^{(022)}$. The steps are quite similar to those from Eq. (40) to Eq. (44). The time derivative of the BH spin in the presence of multiple modes is
@@
$\begin{split}\dot{a}_{*}&=\frac{1}{\tau_{\text{acc}}}\left(\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}\right)\\ &\hskip 28.45274pt-\sum_{l=l_{\text{c}}}^{\infty}\frac{2M_{\text{s}}^{(0ll)}}{M}\left(\frac{l}{M\omega_{0ll}}-2a_{*}\right)\Gamma_{0ll},\end{split}$ (66)
@@
where we have only kept the $\{0,l,l\}$ modes. Other terms have much smaller effect on the BH evolution and could be safely ignored. In the time range from $t_{1}^{(011)}$ to $t_{1}^{(022)}$, only the accretion and the $l=1,2$ modes are important. When the BH follows the $\{0,1,1\}$ Regge curve, the accretion almost balances the superradiance of the $\{0,1,1\}$ mode, resulting in a rather slow change of $a_{*}$. During this time, the $\{0,2,2\}$ mode is unimportant due to its small mass. When the $\{0,2,2\}$ mode is dense enough, it dominates the evolution, causing the rapid drop of $a_{*}$ to the $\{0,2,2\}$ curve close to $t_{1}^{(022)}$. Thus as a LO estimate, one could only consider the contribution of the $\{0,2,2\}$ mode on the right side of Eq. (66)
$\dot{a}_{*}=-\frac{2M_{\text{s}}^{(022)}}{M^{2}\mu}\Gamma_{022}.$ (67)
@@
$M\Gamma_{022}=\frac{64\alpha^{14}}{885735}.$ (68)
The $M_{\text{s}}^{(022)}$ can be calculated by inserting Eqs. (68) and (19) into Eq. (11c) with the GW radiation ignored
$\frac{M_{\text{s}}^{(022)}}{M_{\text{s}}^{(022)}(t_{1}^{(011)})}=\exp\left\{\frac{128{\tau}_{\mathrm{acc}}\mu}{11514555}\left[\alpha^{13}-\left(\alpha_{1}^{(011)}\right)^{13}\right]\right\},$ (69)
which is valid for $t\in[t_{1}^{(011)},t_{1}^{(022)}]$. For the current example, there is $M_{\text{s}}^{(022)}(t_{1}^{(011)})\approx M_{\text{s,0}}^{(022)}$. Then Eq. (67) can be integrated directly from $t_{1}^{(011)}$ to $t_{1}^{(022)}$. Inserting the values of $a_{*\text{c}}^{(0ll)}$ in Eq. (10), one gets
$\begin{split}&\alpha_{1}^{(022)}\approx\Bigg{\{}-\frac{5\cdot 3^{12}}{2^{7}\tau_{\mathrm{acc}}\mu}\times\\ &\left.W_{-1}\left[-\frac{128\tau_{\mathrm{acc}}\mu}{2657205}\left(M_{\mathrm{s0}}^{(022)}\mu e^{-\frac{128\left(\alpha_{1}^{(011)}\right)^{13}\tau_{\mathrm{acc}}\mu}{11514555}}\right)^{\frac{13}{3}}\right]\right\}^{\frac{1}{13}},\end{split}$ (70)
where the term with the exponential integral function is ignored for the same reason discussed below Eq. (44).
The $\alpha_{1}^{(0ll)}$ for larger $l$ can be calculated similarly, which are listed in Eqs. (103). For the current example, after inserting the numbers, one arrives at $\alpha_{1}^{(022)}=0.109$, $\alpha_{1}^{(033)}=0.315$ and $\alpha_{1}^{(044)}=0.639$.
Since $\alpha_{1}^{(011)}-\alpha_{0}\ll 0.1\alpha_{0}$, one could ignore accretion and apply the energy and angular momentum conservations in Eq. (64) for more accurate estimates. Inserting the values, one arrives at $\alpha_{1}^{(011)}=0.097$. The value of $\alpha_{1}^{(022)}$ in Eq. (70) is adjusted accordingly to $0.108$, which satisfies $\alpha_{1}^{(022)}-\alpha_{1}^{(011)}>0.1\alpha_{1}^{(011)}$. Then $\alpha_{1}^{(0ll)}$ for $l\geq 3$ could be calculated using Eqs. (103). Connecting these values on the accretion trajectory as well as the Regge curves, one could get an approximate evolution on the Regge plot, which is shown in Fig. 6(h).

### v2

@@ 4 Multi-mode evolution with accretion
In this section, we study the evolution in the presence of multiple modes. The evolution can be separated into stages with different $l$. In each $l$-stage, the modes with the same azimuthal number $l$ grow the fastest via superradiance. We keep only the dominant ($\{0,l,l\}$) and the subdominant ($\{1,l,l\}$) modes in each $l$-stage for illustration, while the generalization to more modes is straightforward. The superradiance rates of modes with azimuthal numbers larger than $l$ are orders of magnitude smaller and thus can be safely ignored in the $l$-stage. From the experience learned in the previous section, we start with the BH evolution on the Regge plot in Sec. 4.1. Readers not interested in technical details can only read this subsection. Analytical approximations for the important quantities are given in Sec. 4.2. Then the subdominant modes and GW emission flux are discussed in Sec. 4.3 and Sec. 4.4, respectively.
4.1 Evolution on the Regge plot
@@
We first analyze the case when the initial BH spin $a_{*0}$ is greater than the $a_{*\text{c}}^{(0l_{\mathrm{c}}l_{\mathrm{c}})}$ given in Eq. (2.10). If the accretion rate is much larger than the superradiance rate of the $l_{c}$ mode, the BH first follows the accretion trajectory to $\sim a_{*\text{lim}}$ before transiting to the Regge curve with $l=l_{\text{c}}$, see Fig. 6(a). With a decreasing value of the accretion rate, the time cost on the accretion trajectory is shorter, see Fig. 6(b). At some point, the BH spin converges to the $l_{\text{c}}$ Regge curve since $t=0$, see Fig. 6(c). Further decreasing the accretion rate and/or increasing the superradiance rate, the time cost along the $l_{c}$ Regge trajectory is shorter, see Fig. 6(g,h). Until at some point, the accretion effect is negligible during the time when the BH is on the $l_{\text{c}}$ Regge curve. Then the BH does not evolve along the $l_{\text{c}}$ Regge curve before stepping down to the $l_{\text{c}}+1$ Regge trajectory, see Fig. 6(i).
Given the initial conditions, we need to judge which pattern the later evolution has. To separate these cases, one first calculates $\alpha_{0^{\prime}}^{(0l_{\text{c}}l_{\text{c}})}$, which is proportional to the BH mass at the point when its spin reaches the local maximum close to the trajectory curve. The formula for $l_{\text{c}}=1$ is given in Eq. (3.25), and those for larger values of $l_{\text{c}}$ are listed in Eqs. (A.13). When $\mu\tau_{\text{acc}}$ is large, the asymptotic expression of the Lambert $W_{0}$ function could be useful
$\lim_{x\to+\infty}W_{0}(x)=\log x-\log\left(\log x\right).$ (4.2)
@@
In any of the above three cases, the evolution merges onto the $\{0,l_{\text{c}},l_{\text{c}}\}$ Regge curve. With only a single mode as explained in Sec. 3, the BH spin reaches a local minimum close to the Regge curve. There the time of the local minimum is defined as $t_{1}$, and we also define the quantities at this time using the subscript, such as $\alpha_{1}$. In the presence of multiple modes, there is a local minimum of the BH spin at each Regge curve. We similarly define this time at the $\{0,l,l\}$ Regge curve as $t_{1}^{(0ll)}$ and the corresponding quantities using an additional superscript $(0ll)$, such as $\alpha_{1}^{(0ll)}$. To avoid confusion, the mass of a superradiant mode at a certain time is written explicitly. For example $M_{\text{s}}^{(022)}(t_{1}^{(011)})$ denotes the mass of the $\{0,2,2\}$ mode at time $t_{1}^{(011)}$.
The value of $\alpha_{1}^{(0l_{\text{c}}l_{\text{c}})}$ can be obtained using Eqs. (A.14). One could then calculate the value of $\alpha_{1}^{(0ll)}$ with $l>l_{\text{c}}$ using Eqs. (A.20). Then the evolution on the Regge plot can be approximately fixed with the values of $\alpha_{0^{\prime}}$ and $\alpha_{1}^{(0ll)}$. When $\mu\tau_{\text{acc}}$ is large, the following asymptotic expression for the Lambert $W_{-1}$ function is useful
$\lim_{x\to 0^{-}}W_{-1}(x)=\log(-x)-\log\left(-\log(-x)\right).$ (4.3)
@@
${a}_{\mathrm{*c}}^{(0,l+1,l+1)}\left(\alpha^{(0,l+1,l+1)}_{1}\right)^{2}+(l+1)\mu{M}_{\mathrm{s}}^{(0,l+1,l+1)}(t_{1}^{(0ll)})$ $={a}_{\mathrm{*c}}^{(0ll)}\left(\alpha^{(0ll)}_{1}\right)^{2}+l\mu{M}_{\mathrm{s}}^{(0ll)}(t_{1}^{(0ll)}),$ (4.4a)
$\alpha_{1}^{(0,l+1,l+1)}+\mu{M}_{\mathrm{s}}^{(0,1+1,1+1)}(t_{1}^{(0ll)})$ $=\alpha_{1}^{(0ll)}+\mu{M}_{\mathrm{s}}^{(0ll)}(t_{1}^{(0ll)}).$ (4.4b)
The $\alpha_{1}^{(0ll)}$ obtained from these conservation equations agrees better with the numerical results, which can be observed in Fig. 6(c,g-i). Since the conservations give $\alpha_{1}^{(0,l+1,l+1)}\approx 0.9\,\alpha_{1}^{(0ll)}$ [16], we claim Eqs. (4.4) give a more accurate description when $\alpha_{1}^{(0,l+1,l+1)}-\alpha_{1}^{(0ll)}\ll 0.1\,\alpha_{1}^{(0ll)}$.
@@
$a_{*\text{c}}^{(0,l^{\prime}+1,l^{\prime}+1)}<a_{*0}<a_{*\text{c}}^{(0,l^{\prime},l^{\prime})}.$ (4.5)
@@
We define $l_{\text{acc}}$ as the first Regge curve which the evolution follows along. Its value could be larger, smaller, or equal to $l^{\prime}$, depending on the relative size of the accretion rate and the superradiance rates of different modes. We first judge whether $l_{\text{acc}}$ is larger than $l^{\prime}$. After many tests, the best scheme we could find to calculate $l_{\text{acc}}$ is by comparing the $\alpha_{0^{\prime}}^{(0ll)}$ with different $l$. Here $\alpha_{0^{\prime}}^{(0ll)}$ is the mass coupling at the time when the BH spin reaches its local maximum if only accretion and the $\{0,l,l\}$ mode are present, which is given by Eqs. (A.13). Specifically, one calculate the $\alpha_{0^{\prime}}^{(0ll)}$ with increasing value of $l$ from $l_{\text{c}}$ to $l^{\prime}$. If a minimum of $\alpha_{0^{\prime}}^{(0ll)}$ is observed for $l<l^{\prime}$, then the BH spin reaches the local maximum between the $\{0,l-1,l-1\}$ and $\{0,l,l\}$ Regge curves. In this case, we set $l_{\text{acc}}=l$. After identifying $l_{\text{acc}}$, one could reset $t=0$ as the time when the BH reaches the intersection of the accretion trajectory and the $\{0,l_{\text{acc}},l_{\text{acc}}\}$ Regge curve. Then the later evolution can be described in the same way as the $a_{*0}>a_{*\text{c}}^{(0l_{\text{c}}l_{\text{c}})}$ case with $l_{\text{c}}$ replaced by $l_{\text{acc}}$.
@@
The evolutions with $a_{*0}=0$ are presented in Fig. 6, panels (d-f) and panels (j-l), where other initial parameters are the same as panels (a-c) and panels (g-i), respectively. In this case, the $l^{\prime}$ is infinity. Analytical estimates of the points where the evolution leaves a trajectory and merges into the next are given in Sec. 4.2. For various initial conditions and accretion rates, the analytical approximations agree quite well with the numerical results.
4.2 Evolution of the BH and dominate modes
@@
The BH evolution is mainly determined by the dominant mode and/or the accretion in each $l$-stage. In this part, we focus on the dominant modes, while leaving the discussion about the subdominant modes in Sec. 4.3. To help the analysis, we present the numerical results for a specific multi-mode scenario in Fig. 7. The initial parameters are $M_{0}=10^{4}M_{\odot}$, $\alpha_{0}=0.1$, $M_{\text{s,0}}^{(nll)}=10^{-5}M_{0}$ and $f_{\text{Edd}}=10^{-3}$. The value of $f_{\text{Edd}}$ is chosen such that the accretion time scale ${\tau}_{\text{acc}}$ is at the same order of $\tau_{\text{SR}}^{(022)}$, which is the the superradiance time scale of $(0,2,2)$ mode. This time scale is much larger than $\tau_{\text{SR}}^{(011)}$ but much smaller than $\tau_{\text{SR}}^{(033)}$.
@@
Two initial BH spins $a_{*0}=0.7$ and $a_{*0}=0$ are considered. From Eq. (4.1), it is easy to see $l_{\text{c}}=1$. Then $a_{*\text{c}}^{(011)}\approx 0.384$ is calculated from Eq. (2.10) using the value of $\alpha_{0}$.
4.2.1 $a_{*0}=0.7$
@@
This case corresponds to $a_{*0}>a_{*\text{c}}^{(0l_{\mathrm{c}}l_{\mathrm{c}})}$. We first identify the evolution of the BH on the Regge plot. From Eq. (2.18), one could obtain $\tau_{\text{acc}}\mu=3.2\times 10^{18}$. We then apply Eq. (3.25) and get $\alpha_{0^{\prime}}=\alpha_{0}-8\times 10^{-10}$, which is less than $\alpha_{0}$. Thus the BH does not evolve along the accretion trajectory at all, corresponding to the third scenario in the discussion below Eq. (4.2).
@@
The $\alpha_{1}^{(011)}$ can be calculated using Eq. (3.11). Since the BH does not evolve along the accretion trajectory, the $M_{\text{s},0^{\prime\prime}}$ $a_{*0^{\prime\prime}}$, and $\alpha_{0^{\prime\prime}}$ should be replaced by $M_{\text{s,0}}$, $a_{*0}$, and $\alpha_{0}$, respectively. Using the above parameters, one obtains $\alpha_{1}^{(011)}=\alpha_{0}+7\times 10^{-10}$. The small values of $\alpha_{0^{\prime}}-\alpha_{0}$ and $\alpha_{1}^{(011)}-\alpha_{0}$ indicate the evolution merges onto the $\{0,1,1\}$ Regge curve almost at $t=0$.
@@
We continue to calculate $\alpha_{1}^{(022)}$. The steps are quite similar to those from Eq. (3.7) to Eq. (3.11). The time derivative of the BH spin in the presence of multiple modes is
@@
$\dot{a}_{*}$ $=\frac{1}{\tau_{\text{acc}}}\left(\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}\right)-\sum_{l=l_{\text{c}}}^{\infty}\frac{2M_{\text{s}}^{(0ll)}}{M}\left(\frac{l}{M\omega_{0ll}}-2a_{*}\right)\Gamma_{0ll},$ (4.6)
@@
where we have only kept the $\{0,l,l\}$ modes. Other terms have much smaller effect on the BH evolution and could be safely ignored. In the time range from $t_{1}^{(011)}$ to $t_{1}^{(022)}$, only the accretion and the $l=1,2$ modes are important. When the BH follows the $\{0,1,1\}$ Regge curve, the accretion almost balances the superradiance of the $\{0,1,1\}$ mode, resulting in a rather slow change of $a_{*}$. During this time, the $\{0,2,2\}$ mode is unimportant due to its small mass. When the $\{0,2,2\}$ mode is dense enough, it dominates the evolution, causing the rapid drop of $a_{*}$ to the $\{0,2,2\}$ curve close to $t_{1}^{(022)}$. Thus as a LO estimate, one could only consider the contribution of the $\{0,2,2\}$ mode on the right side of Eq. (4.6)
$\dot{a}_{*}=-\frac{2M_{\text{s}}^{(022)}}{M^{2}\mu}\Gamma_{022}.$ (4.7)
@@
$M\Gamma_{022}=\frac{64\alpha^{14}}{885735}.$ (4.8)
The $M_{\text{s}}^{(022)}$ can be calculated by inserting Eqs. (4.8) and (2.19) into Eq. (2.11c) with the GW radiation ignored
$\frac{M_{\text{s}}^{(022)}}{M_{\text{s}}^{(022)}(t_{1}^{(011)})}=\exp\left\{\frac{128{\tau}_{\mathrm{acc}}\mu}{11514555}\left[\alpha^{13}-\left(\alpha_{1}^{(011)}\right)^{13}\right]\right\},$ (4.9)
which is valid for $t\in[t_{1}^{(011)},t_{1}^{(022)}]$. For the current example, there is $M_{\text{s}}^{(022)}(t_{1}^{(011)})\approx M_{\text{s,0}}^{(022)}$. Then Eq. (4.7) can be integrated directly from $t_{1}^{(011)}$ to $t_{1}^{(022)}$. Inserting the values of $a_{*\text{c}}^{(0ll)}$ in Eq. (2.10), one gets
$\alpha_{1}^{(022)}\approx\left\{-\frac{5\cdot 3^{12}}{2^{7}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{128\tau_{\mathrm{acc}}\mu}{2657205}\left(M_{\mathrm{s0}}^{(022)}\mu e^{-\frac{128\left(\alpha_{1}^{(011)}\right)^{13}\tau_{\mathrm{acc}}\mu}{11514555}}\right)^{\frac{13}{3}}\right]\right\}^{\frac{1}{13}},$ (4.10)
where the term with the exponential integral function is ignored for the same reason discussed below Eq. (3.11).
The $\alpha_{1}^{(0ll)}$ for larger $l$ can be calculated similarly, which are listed in Eqs. (A.20). For the current example, after inserting the numbers, one arrives at $\alpha_{1}^{(022)}=0.109$, $\alpha_{1}^{(033)}=0.315$ and $\alpha_{1}^{(044)}=0.639$.
Since $\alpha_{1}^{(011)}-\alpha_{0}\ll 0.1\alpha_{0}$, one could ignore accretion and apply the energy and angular momentum conservations in Eq. (4.4) for more accurate estimates. Inserting the values, one arrives at $\alpha_{1}^{(011)}=0.097$. The value of $\alpha_{1}^{(022)}$ in Eq. (4.10) is adjusted accordingly to $0.108$, which satisfies $\alpha_{1}^{(022)}-\alpha_{1}^{(011)}>0.1\alpha_{1}^{(011)}$. Then $\alpha_{1}^{(0ll)}$ for $l\geq 3$ could be calculated using Eqs. (A.20). Connecting these values on the accretion trajectory as well as the Regge curves, one could get an approximate evolution on the Regge plot, which is shown in Fig. 6(h).